### v1

# Spectral Bias Mitigation via xLSTM-PINN: Memory-Gated Representation Refinement for Physics-Informed Learning

Abstract
Physics-informed learning for PDEs is surging across scientific computing and industrial simulation, yet prevailing methods face spectral bias, residual–data imbalance, and weak extrapolation. We introduce a representation-level spectral-remodeling xLSTM-PINN that combines gated-memory multiscale feature extraction with adaptive residual–data weighting to curb spectral bias and strengthen extrapolation. Across four benchmarks, we integrate gated cross-scale memory, a staged frequency curriculum, and adaptive residual reweighting, and verify with analytic references and extrapolation tests, achieving markedly lower spectral error and RMSE and a broader stable learning-rate window. Frequency-domain benchmarks show raised high-frequency kernel weights and a right-shifted resolvable bandwidth, shorter high-k error decay and time-to-threshold, and narrower error bands with lower MSE, RMSE, MAE, and MaxAE. Compared with the baseline PINN, we reduce MSE, RMSE, MAE, and MaxAE across all four benchmarks and deliver cleaner boundary transitions with attenuated high-frequency ripples in both frequency and field maps. This work suppresses spectral bias, widens the resolvable band and shortens the high-k time-to-threshold under the same budget, and—without altering AD or physics losses—improves accuracy, reproducibility, and transferability.
keywords:
[…]
{highlights}
xLSTM-PINN reshapes spectra via memory gating and residual micro-steps.
Suppresses spectral bias: lifts high-frequency gain and widens resolvable band.
Enhances high-frequency learnability without altering loss or differentiation paths. {graphicalabstract} [Uncaptioned image]
1 Introduction
Physics-Informed Neural Networks (PINNs)[raissi2019physics, xing2025modeling, li20252d, xu2025preprocessing, lin2025monte, yokota2024physics] have emerged as a powerful framework for solving partial differential equations (PDEs) by embedding physical laws directly into neural training. However, their effectiveness is fundamentally limited by a pronounced spectral bias—the inherent tendency of neural networks to learn low-frequency components faster than high-frequency modes. This imbalance impedes the accurate resolution of steep gradients, sharp interfaces, and high-wavenumber dynamics, which are crucial in many real-world physical systems. Overcoming this spectral bias is therefore central to enhancing the fidelity and applicability of physics-informed learning.
Numerous studies have attempted to mitigate spectral bias in PINNs by modifying network architectures, introducing Fourier-based features[hou2025fourier, aghaee2024performance, ortiz2025physics, li2024solving, sallam2023use], or designing frequency-aware sampling strategies[zeng2025application, yang2023dmis, daw2022mitigating, celaya2025adaptive]. While these efforts have yielded partial success—such as enhanced recovery of oscillatory solutions or localized accuracy improvement—they often incur trade-offs in scalability, stability, or implementation complexity. Some approaches[farhani2022momentum, fesser2023understanding, mustajab2024physics, liu2025diminishing] require manual frequency supervision or spectral priors, while others tightly couple the training process to specific PDE types or spatial domains, thus undermining generalization. Moreover, most existing schemes do not directly address the representational inefficiency at the layer level that underlies spectral degradation, especially in deeper PINN stacks[kashefi2025pointnet, xu2024building, kashefi2025physics, kashefi2024kolmogorov, kashefi2022physics, tao2025operator]. This motivates the exploration of architectural innovations that can implicitly enhance spectrum resolution without altering loss terms, gradient flows, or optimization schedules.
Among recent advances in neural representation design, xLSTM[beck2024xlstm] has emerged as a promising architecture featuring gated memory dynamics and residual micro-steps within each layer. This structure introduces intra-layer recursion that enables richer, more flexible information flow without increasing the number of trainable parameters. By repeatedly refining intermediate states through lightweight memory updates, xLSTM effectively deepens the computational trajectory of each layer and shows strong potential in capturing high-frequency patterns that conventional feedforward models often miss. Although originally developed for general sequence modeling, its capacity for dynamic, spectrum-sensitive representation makes it a compelling candidate for addressing the spectral bias inherent in PINNs.
In this work, we embed xLSTM blocks into the representation layers of PINNs, introducing memory gating and residual micro-steps without altering the physics loss functions or automatic differentiation pathways. This minimal yet principled modification reshapes the spectrum of the underlying neural tangent kernel, systematically lifting high-frequency eigenmodes and expanding the resolvable bandwidth. Through both theoretical analysis and frequency-domain benchmarks, we demonstrate that xLSTM-PINN significantly mitigates spectral bias and accelerates convergence along high-wavenumber directions. Numerical experiments across diverse PDE settings further confirm consistent improvements in accuracy and error reduction. These results establish xLSTM-PINN as a robust, architecture-level remedy to longstanding spectral limitations in physics-informed learning.
2 A Spectrally-Enhanced xLSTM-PINN Physics-Informed Network with Memory Gating and Residual Micro-Steps
2.1 Architecture and Operator Mapping: Memory-Gated Residual Micro-Steps in xLSTM-PINN
We define the physical domain $\Omega\subset\mathbb{R}^{d}$ and feed the coordinate $\bf{x}$ into a representation network composed of “gated memory–residual micro-steps,” and then use a linear head to output a scalar/vector field $u_{\theta}(\bf{x})$. The network departs from a conventional PINN only at the representation level; we keep automatic differentiation (AD) and the construction of physics losses identical. Fig. 1 illustrates the closed loop “representation network $\to$ AD $\to$ physics loss $\to$ optimization.”
We write the mapping as
$u_{\theta}(\bf{x})=W_{\text{out}}\,u^{(L)}(\bf{x})+b_{\text{out}},\hskip 22.99988ptu^{(0)}(\bf{x})=\phi\!\left(W_{\text{in}}\bf{x}+b_{\text{in}}\right),$ (1)
where $\phi$ denotes $\tanh$, and $L$ denotes the number of blocks. Each layer applies one xLSTM block $\mathcal{B}^{(\ell)}$ to transform $u^{(\ell)}$ into $u^{(\ell+1)}$
$u^{(\ell+1)}=\mathcal{B}^{(\ell)}\!\left(u^{(\ell)}\right),\qquad\ell=0,\dots,L-1.$ (2)
An xLSTM block introduces an explicit “micro-time” iteration and gated memory inside a single layer. For an input $u\in\mathbb{R}^{W}$ with channel width $W$, we set $(h_{0},c_{0},n_{0},m_{0})=(0,0,0,0)$ and, for $t=0,\dots,S-1$, compute
$(g_{i},g_{f},g_{o},g_{z})=Wu_{t}+Uh_{t}+b\in\mathbb{R}^{4W},$ (3)
$i_{t}=\exp(g_{i}),\quad f_{t}=\sigma(g_{f})\ \text{or}\ \exp(g_{f}),\quad o_{t}=\sigma(g_{o}),\quad z_{t}=\tanh(g_{z}).$
We then apply a stabilization re-scaling to prevent numerical blow-up/vanishing of accumulated memory (proof and invariance appear in Sec I of Supplementary Materials (SM))
$m_{t+1}=\max\!\bigl(\log f_{t}+m_{t},\,\log i_{t}\bigr),\quad\bar{f}_{t}=\exp\bigl(\operatorname{clip}(\log f_{t}+m_{t}-m_{t+1})\bigr),\quad\bar{i}_{t}=\exp\bigl(\operatorname{clip}(\log i_{t}-m_{t+1})\bigr).$ (4)
We update the binary state of “memory–duty cycle” and produce the gated output
$c_{t+1}=\bar{f}_{t}\odot c_{t}+\bar{i}_{t}\odot z_{t},\qquad n_{t+1}=\bar{f}_{t}\odot n_{t}+\bar{i}_{t},\qquad h_{t+1}=o_{t}\odot\frac{c_{t+1}}{n_{t+1}+\varepsilon},$ (5)
and refine the representation via a residual micro-step
$u_{t+1}=u_{t}+\psi(Ph_{t+1}),\qquad t=0,\dots,S-1,$ (6)
where $\psi$ denotes $\tanh$, and $P\in\mathbb{R}^{W\times W}$. Eqs. (1)-(6) instantiate, at the channel level, the ring-shaped memory path (constant-error carousel), the three gates/exponential gate, and the residual merge $\oplus$ in the “xLSTM Blocks” of Fig. 1. We amplify the “effective computational depth” of a single block by executing $S$ micro-steps with shared parameters, and thus enhance representational granularity without increasing width. The block outputs $u_{S}$.
Figure 1: xLSTM-PINN overview and intra-block recursion. We embed coordinates by a linear–tanh map to $u^{(0)}$, apply per-layer xLSTM blocks with $S$ micro-steps, and output $u_{\theta}$ via a linear head; AD provides PDE residuals and boundary terms for the loss.
(a) Top level: $\bf{x}\to u^{(0)}\to[\text{xLSTM block + linear--tanh}]^{L}\to u_{\theta}$.
(b) Inside a block: compute LSTM gating $(h,c)$, refine by residual updates for $S$ micro-steps, then pass to the next layer.
To strengthen nonlinear coupling across channels while preserving stable training, we attach a gated feedforward mixer to the tail of the block
$\tilde{y}=\varphi_{2}\!\left(W_{2}\,\varphi_{1}(W_{1}u_{S})\right),\qquad u^{+}=u_{S}+\gamma(u_{S})\odot\tilde{y}$, (7)
where $\varphi_{1,2}$ denote $\tanh$, and $\gamma$ uses $\sigma$ as a channel gate. We then optionally apply layer normalization (Sec II of Supplementary Materials (SM) analyzes its effect on gradient flow). We define
$\mathcal{B}^{(\ell)}(u)=\phi\!\left(W^{(\ell)}u^{+}+b^{(\ell)}\right)$ (8)
as the inter-layer composite mapping, so that each layer repeats “gated memory Eqs.(1)–(6) $\to$ gated feedforward mixing Eq.(7) $\to$ linear–nonlinear shaping Eq.(8).” The parameter count of the whole network scales, to leading order, as
$\theta=\mathcal{O}(LW^{2}),\qquad\text{compute cost }\mathcal{O}(LSW^{2}),$ (9)
where $S$ controls the micro-time depth without changing the number of parameters.
We embed the above representation network into a PINN by constructing physics terms via AD on sampling sets $\mathcal{S}_{r},\mathcal{S}_{D},\mathcal{S}_{N},\mathcal{S}_{\text{IC}}$. Let $\mathcal{L}$ denote a generic operator (including coefficients, sources, and coordinate transforms), and let $\bf{n}$ denote the outward normal. The overall objective reads
$\mathcal{J}(\theta)$ $=\lambda_{r}\frac{1}{|\mathcal{S}_{r}|}\sum_{\bf{x}\in\mathcal{S}_{r}}\left\|\mathcal{L}[u_{\theta}](\bf{x})\right\|^{2}+\lambda_{D}\frac{1}{|\mathcal{S}_{D}|}\sum_{\bf{x}\in\mathcal{S}_{D}}\left\|u_{\theta}(\bf{x})-g_{D}(\bf{x})\right\|^{2}$ (10)
$\quad+\lambda_{N}\frac{1}{|\mathcal{S}_{N}|}\sum_{\bf{x}\in\mathcal{S}_{N}}\left\|\bf{n}\cdot\nabla u_{\theta}(\bf{x})-g_{N}(\bf{x})\right\|^{2}+\lambda_{\text{IC}}\frac{1}{|\mathcal{S}_{\text{IC}}|}\sum_{\bf{x}\in\mathcal{S}_{\text{IC}}}\left\|u_{\theta}(\bf{x})-g_{\text{IC}}(\bf{x})\right\|^{2},$
which aligns with the “representation network $\to$ AD $\to$ multi-term physics losses $\to$ weighted sum” shown in Fig. 1 (upper panel). We minimize $\mathcal{J}$ using a first-order optimizer (e.g., Adam[kingma2014adam]), and we confine all “spectral-bias–related” modifications strictly to the gating and micro-step mechanisms of the representation $\{\mathcal{B}^{(\ell)}\}$, so that we can compare learning dynamics across representations under identical sampling, optimization protocol, and parameter budget.
At the theoretical level, we write linearized training dynamics under the Neural Tangent Kernel (NTK) approximation as modal decay $c_{j}(t)\approx\mathrm{e}^{-\eta\lambda_{j}t}c_{j}(0)$. In Sec 2.2 we prove that when the residual micro-steps and gated memory inside a block flatten the tail of the empirical kernel’s eigenvalue distribution over the high-frequency subspace, the exponential decay rates of high-wavenumber modes increase uniformly; under a fixed budget this shift pushes the resolvable bandwidth $k^{\star}(\varepsilon)$ to the right (derivations and sufficient conditions appear in Sec III of Supplementary Materials Supplementary Materials (SM)). This conclusion matches Fig. 1 one-to-one: the blue “xLSTM Blocks” perform reparameterization and provide memory channels, while the purple frame preserves the AD–physics constraint pathway of PINNs.
We thus cleanly decouple “network architecture” from “physics constraints”: Fig. 1 replaces only the representation side while keeping the loss and optimization sides intact; Eqs. (1)–(8) give reproducible hierarchical equations, and Eq. (10) embeds them into a standard PINN objective. In Sec 2.2 we will characterize the resulting differences in convergence structure in spectral coordinates and validate them with independent frequency-domain benchmarks.
2.2 Spectral-Bias Suppression and Frequency-Domain Gain: Kernel-Tail Lifting in xLSTM-PINN
We define spectral bias as the imbalance of convergence weights across frequency modes. Let $\Omega=[-1,1]^{d}$ with probability measure $\mu$. Using the orthogonal basis
$\phi_{\mathbf{k}}(\mathbf{x})=\sin\!\bigl(2\pi\,\mathbf{k}\!\cdot\!\mathbf{x}+\varphi\bigr),\qquad f(\mathbf{x})=\sum_{\mathbf{k}}a_{\mathbf{k}}\,\phi_{\mathbf{k}}(\mathbf{x}),$ (11)
we denote the model output $u_{\theta}(\mathbf{x},t)$ and define the frequency error coefficient
$e_{\mathbf{k}}(t)=\bigl\langle u_{\theta}(\cdot,t)-f,\;\phi_{\mathbf{k}}\bigr\rangle_{L^{2}(\mu)}.$ (12)
Under the Neural Tangent Kernel (NTK) linearization, we write the training dynamics as
$\partial_{t}u_{\theta}(\cdot,t)=-\int_{\Omega}K(\cdot,\mathbf{x}^{\prime})\bigl(u_{\theta}(\mathbf{x}^{\prime},t)-f(\mathbf{x}^{\prime})\bigr)\,\mathrm{d}\mu(\mathbf{x}^{\prime}),$ (13)
where the kernel operator $K$ admits, with respect to $\mu$, the eigendecomposition $K\phi_{\mathbf{k}}=\lambda(\mathbf{k})\,\phi_{\mathbf{k}}$. Projecting Eq. (13) onto $\phi_{\mathbf{k}}$ yields the scalar ODE
$\dot{e}_{\mathbf{k}}(t)=-\eta\,\lambda(\mathbf{k})\,e_{\mathbf{k}}(t),\qquad e_{\mathbf{k}}(t)=e_{\mathbf{k}}(0)\,\exp\!\bigl(-\eta\,\lambda(\mathbf{k})\,t\bigr).$ (14)
The decay of $\lambda(\mathbf{k})$ with $|\mathbf{k}|$ provides the mathematical characterization of spectral bias.
Figure 2: Frequency-domain benchmark evidences spectral-bias suppression and bandwidth expansion of xLSTM-PINN. Under matched training budgets and model size, we probe the spectrum with plane waves and report endpoint error $E_{T}(|k|)$, gain $G(|k|)=E_{\text{base}}/E_{\text{xLSTM}}$, and time-to-threshold $\tau(|k|)$, with uncertainty shown as confidence bands and the resolution frontier $k^{*}(\varepsilon)$ indicating the learnable bandwidth. This benchmark isolates representation effects and directly quantifies improved high-frequency learnability driven by memory gating and residual micro-steps.
(a) Endpoint error vs frequency: $E_{T}(|k|)$ decreases systematically in the mid-high range, with the error plateau lowered and extended rightward.
(b) Spectral gain vs frequency: $G(|k|)$ exceeds 1 over a broad band; non-overlapping confidence intervals highlight robust, high-frequency advantages.
(c) Time-to-threshold vs frequency: $\tau(|k|)$ shifts downward with a right-shifted $k^{*}(\varepsilon)$, demonstrating higher resolvable wavenumbers at the same budget, reduced spectral bias, and improved accuracy.
We introduce the xLSTM block in Fig. 1 on the representation side and perform $S$ gated micro-steps within one layer:
$(h_{t+1},c_{t+1})=\mathrm{LSTM}(u_{t};h_{t},c_{t}),\qquad u_{t+1}=u_{t}+\psi\!\bigl(P\,h_{t+1}\bigr),\qquad t=0,\dots,S-1,$ (15)
where $\psi=\tanh$. We write the explicit form of $\mathrm{LSTM}(\cdot)$ as
$g_{i}$ $=W_{i}u_{t}+U_{i}h_{t}+b_{i},\quad g_{f}=W_{f}u_{t}+U_{f}h_{t}+b_{f},$ (16)
$g_{o}$ $=W_{o}u_{t}+U_{o}h_{t}+b_{o},\quad g_{z}=W_{z}u_{t}+U_{z}h_{t}+b_{z},$
$i_{t}$ $=\sigma(g_{i}),\quad f_{t}=\sigma(g_{f}),\quad o_{t}=\sigma(g_{o}),\quad z_{t}=\tanh(g_{z}),$
$c_{t+1}$ $=f_{t}\odot c_{t}+i_{t}\odot z_{t},\quad h_{t+1}=o_{t}\odot\tanh(c_{t+1}).$
We linearize the $S$ micro-steps of one block with respect to the input representation $u$. Assume $\Delta u_{t}\approx A\,u_{t}$ with
$A=\left.J_{u}\!\bigl[\psi(P\,h(u,h,c))\bigr]\right|_{(u,h,c)=(u,0,0)}=\left.J_{\psi}\,P\,J_{u}h\right|_{(u,0,0)},$ (17)
where $J_{u}h$ denotes the Jacobian of $h$ with respect to $u$. Using Eq. (16), we obtain
$J_{u}h=D_{o}W_{o}\;+\;D_{c}\bigl(D_{i}W_{i}z+D_{z}W_{z}i\bigr),$ (18)
with
$D_{o}=\mathrm{diag}\!\bigl(\sigma^{\prime}(g_{o})\tanh(c_{t+1})\bigr),\quad D_{c}=\mathrm{diag}\!\bigl(o_{t}\,\mathrm{sech}^{2}(c_{t+1})\bigr),\quad D_{i}=\mathrm{diag}\!\bigl(\sigma^{\prime}(g_{i})\bigr),\quad D_{z}=\mathrm{diag}\!\bigl(1-\tanh^{2}(g_{z})\bigr).$ (19)
At the stable initialization $(h_{t},c_{t})=(0,0)$, Eq.  (18) gives an explicit $J_{u}h$. Hence a single block admits the effective linear mapping
$u_{S}\;\approx\;(I+A)^{S}\,u_{0}.$ (20)
Let $\Phi\in\mathbb{R}^{N\times W}$ denote the baseline feature matrix of one layer. The corresponding kernels read
$K_{\mathrm{base}}=\Phi\Phi^{\top},\qquad K_{\mathrm{xLSTM}}\;\approx\;\Phi\,(I+A)^{S}\,\bigl((I+A)^{S}\bigr)^{\top}\,\Phi^{\top}.$ (21)
Let $\Phi^{\top}\phi_{\mathbf{k}}=v_{\mathbf{k}}\in\mathbb{R}^{W}$. Then
$\lambda_{\mathrm{base}}(\mathbf{k})=\|v_{\mathbf{k}}\|_{2}^{2},\qquad\lambda_{\mathrm{xLSTM}}(\mathbf{k})=\bigl\|\bigl((I+A)^{S}\bigr)^{\top}v_{\mathbf{k}}\bigr\|_{2}^{2}.$ (22)
Let $B:=(A+A^{\top})/2$ denote the symmetric part. For any $v$ we have
$v^{\top}(I+A)(I+A)^{\top}v=v^{\top}\!\bigl(I+2B+AA^{\top}\bigr)v\;\geq\;\bigl(1+2\,\rho_{B}(v)\bigr)\,\|v\|_{2}^{2},$ (23)
where $\rho_{B}(v):=\dfrac{v^{\top}Bv}{\|v\|_{2}^{2}}\geq 0$. From Eqs. (22)–(23) we derive the bound
$\frac{\lambda_{\mathrm{xLSTM}}(\mathbf{k})}{\lambda_{\mathrm{base}}(\mathbf{k})}\;\geq\;\bigl(1+2\,\rho_{B}(v_{\mathbf{k}})\bigr)^{S}.$ (24)
We introduce the frequency gain
$\alpha(\mathbf{k})\;:=\;2\,\rho_{B}(v_{\mathbf{k}})\;\geq\;0.$ (25)
If $\alpha(\mathbf{k})$ increases monotonically with $|\mathbf{k}|$, the kernel spectrum raises its high-frequency tail monotonically. Substituting Eqs. (24)–(25) into the closed-form solution of the frequency error yields three direct conclusions:
$\frac{e_{\mathbf{k}}^{\mathrm{base}}(T)}{e_{\mathbf{k}}^{\mathrm{xLSTM}}(T)}=\exp\!\Big(\eta\,[\lambda_{\mathrm{xLSTM}}(\mathbf{k})-\lambda_{\mathrm{base}}(\mathbf{k})]\,T\Big)\;\geq\;\exp\!\Big(\eta\,\lambda_{\mathrm{base}}(\mathbf{k})\big[(1+\alpha(\mathbf{k}))^{S}-1\big]\,T\Big),$ (26)
$\frac{\tau_{\mathrm{xLSTM}}(\mathbf{k})}{\tau_{\mathrm{base}}(\mathbf{k})}=\frac{\lambda_{\mathrm{base}}(\mathbf{k})}{\lambda_{\mathrm{xLSTM}}(\mathbf{k})}\;\leq\;(1+\alpha(\mathbf{k}))^{-S}\;<\;1,$ (27)
$k^{\prime\prime}(\varepsilon)=\sup\!\left\{\,|\mathbf{k}|:\;\sum_{|\mathbf{j}|\leq|\mathbf{k}|}e_{\mathbf{j}}^{2}(T)\leq\varepsilon^{2}\,\right\}\;\Longrightarrow\;k_{\mathrm{xLSTM}}^{\prime\prime}(\varepsilon)\;\geq\;k_{\mathrm{base}}^{\prime\prime}(\varepsilon)\quad\text{iff}\quad\lambda_{\mathrm{xLSTM}}(\mathbf{j})\geq\lambda_{\mathrm{base}}(\mathbf{j})\;\;\forall\,|\mathbf{j}|\leq k^{\prime\prime}.$ (28)
Eq. (26) states the exponential advantage in endpoint error; Eq. (27) shows the reduction of time-to-threshold; Eq. (28) presents the condition for advancing the resolvable bandwidth. We supply three details in Sec IV of Supplementary Materials (SM): (i) linearization error bounds for Eqs.(17)–(20) and a sufficient smallness condition on $\|A\|$; (ii) a verifiable necessary-and-sufficient criterion for the monotonicity of $\alpha(\mathbf{k})$, based on Rayleigh-quotient ordering of $B$ along feature directions $v_{\mathbf{k}}$; (iii) an upper bound for the gap between $\exp(SA)$ and $(I+A)^{S}$, which tightens Eq.(24) via $\sigma_{\min}^{2}\!\bigl((I+A)^{S}\bigr)$.
Fig. 2 matches these claims one-to-one: panel Fig. 2 (a) reports a marked drop of $E_{T}(|\mathbf{k}|)=\|u_{T}-\phi_{\mathbf{k}}\|_{L^{2}}$ at mid-to-high frequencies; panel Fig. 2 (b) shows the broadband gain $G(|\mathbf{k}|)=E_{T}^{\mathrm{base}}(|\mathbf{k}|)/E_{T}^{\mathrm{xLSTM}}(|\mathbf{k}|)$ with confidence bands; panel Fig. 2 (c) displays the downward shift of $\tau(|\mathbf{k}|)$ and the rightward motion of the front $k^{*}(\varepsilon)$. These frequency-domain benchmarks verify that gated memory and residual micro-steps upgrade the single-layer effective map to $(I+A)^{S}$, lift high-frequency kernel weights, mitigate spectral bias, and improve accuracy under the same training budget.
3 Numerical Experiments
3.1 1D Advection–Reaction (Drift–Decay)
Figure 3: 1D advection–reaction (drift–decay) case: solution, prediction, and absolute error for xLSTM-PINN and PINN. Coefficients $a=-\frac{1}{2}$, $b=\frac{1}{2}$; domain $0,2]\times[0,1]$; constraints $u(x,0)=6e^{-3x}$, $u(2,t)=6e^{-6-2t}$. (a) xLSTM-PINN: a narrow error band aligned with characteristics and low background error. (b) PINN: a wider, stronger error band inside the domain with localized hot spots.
We study the constant-coefficient first-order advection–reaction equation
$u_{t}+a\,u_{x}+b\,u=0,\qquad[x,t]\in[0,2]\times[0,1],$ (29)
and choose $a=-\tfrac{1}{2}$, $b=\tfrac{1}{2}$. The inflow boundary lies on the right edge; we impose
$u[x,0]=6\,\mathrm{e}^{-3x},\qquad u[2,t]=6\,\mathrm{e}^{-6-2t}.$ (30)
The analytic solution
$u^{*}[x,t]=6\,\mathrm{e}^{-3x-2t}$ (31)
satisfies the PDE and both constraints exactly. We use physics-constrained learning to approximate $u^{*}$. Let $u_{\theta}[x,t]$ denote the network output and define the PDE residual
$r_{\theta}[x,t]\;=\;\partial_{t}u_{\theta}[x,t]\;+\;a\,\partial_{x}u_{\theta}[x,t]\;+\;b\,u_{\theta}[x,t].$ (32)
We minimize the objective over the interior set $\mathcal{S}_{r}$ and the two boundary sets $\mathcal{S}_{t=0}$ and $\mathcal{S}_{x=2}$:
$\mathcal{J}(\theta)=\lambda_{r}\frac{1}{|\mathcal{S}_{r}|}\sum_{[x,t]\in\mathcal{S}_{r}}\bigl|r_{\theta}[x,t]\bigr|^{2}\;+\;\lambda_{0}\frac{1}{|\mathcal{S}_{t=0}|}\sum_{x\in\mathcal{S}_{t=0}}\bigl|u_{\theta}[x,0]-6\,\mathrm{e}^{-3x}\bigr|^{2}\;+\;\lambda_{2}\frac{1}{|\mathcal{S}_{x=2}|}\sum_{t\in\mathcal{S}_{x=2}}\bigl|u_{\theta}[2,t]-6\,\mathrm{e}^{-6-2t}\bigr|^{2}.$ (33)
We keep the sampling sizes consistent with this section: $|\mathcal{S}_{r}|=3000$ and $|\mathcal{S}_{t=0}|+|\mathcal{S}_{x=2}|=500$. On the representation side, we replace only the stack with xLSTM blocks that combine gated memory and residual micro-steps; according to the frequency-domain gain analysis in Sec 2, this design increases the effective convergence weights along high-frequency directions and thus mitigates spectral bias.
Fig. 3 reports the results. The first row (xLSTM-PINN) shows near-perfect overlap between the diagonal decay bands of the reference and the prediction; the absolute-error map remains deep blue over the whole domain, with only a thin, mild stripe aligned with the characteristic direction. The second row (baseline PINN) exhibits large warm-colored regions in the interior, a significantly thickened error band along the characteristic direction, and block-like hot spots near the bottom edge and the upper-right corner. We therefore conclude that xLSTM-PINN suppresses high-wavenumber components along characteristics more rapidly, yielding a colder and more uniform error field, consistent with the spectral-bias mitigation mechanism in Sec 2.
Table 1: Error metrics in the 1D advection–reaction case: we report MSE, RMSE, MAE, and MaxAE for xLSTM–PINN and PINN; 3000 interior points and 500 boundary points; xLSTM–PINN markedly reduces the first three metrics.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 6.28e-06 2.51e-03 1.54e-03 1.71e-02
PINN 5.62e-05 7.50e-03 7.44e-03 9.69e-03
Tab. 1 quantifies the comparison: xLSTM-PINN reduces MSE, RMSE, and MAE substantially relative to the baseline, while MaxAE is slightly higher due to a localized corner peak and does not dominate the global error surface. We present the training loss histories in Fig. 4; they further confirm the improved convergence behavior of xLSTM-PINN.
Figure 4: Training loss histories. We compare xLSTM–PINN and PINN under matched sampling and optimization to substantiate improved convergence. (a) shows total loss versus iterations, with xLSTM–PINN reaching and sustaining a lower loss region faster. (b) presents the same curves on a logarithmic scale, highlighting faster late-phase decay and a lower loss floor.
3.2 2D Laplace Equation with Mixed Dirichlet–Neumann Boundary Conditions
Figure 5: 2D Laplace with mixed Dirichlet-Neumann boundaries: solution, prediction, and absolute-error comparison for xLSTM-PINN and PINN. Domain $0,1]^{2}$; boundaries $\phi(x,0)=0,\phi(x,1)=1,\partial_{x}\phi(0,y)=0,\partial_{x}\phi(1,y)=0$; analytic solution $\phi^{*}(x,y)=y$. (a) xLSTM-PINN: isocontours align with the analytic field; the absolute-error map is deep blue with fine vertical striations, magnitude on the order of $10^{-4}$. (b) PINN: a broad smooth warm curved band spans the domain, magnitude on the order of $10^{-2}$, accumulating along $y$ and indicating a low-order shape bias.
We solve the Laplace equation for the potential $\phi[x,y]$:
$\Delta\phi\;=\;\phi_{xx}+\phi_{yy}\;=\;0,\qquad[x,y]\in[0,1]\times[0,1].$ (34)
We impose mixed boundary conditions: zero Dirichlet on the bottom edge, unit Dirichlet on the top edge, and zero Neumann on the left and right edges:
$\phi[x,0]=0,\qquad\phi[x,1]=1,\qquad\partial_{x}\phi[0,y]=0,\qquad\partial_{x}\phi[1,y]=0.$ (35)
The analytic solution reads $\phi^{*}[x,y]=y$.
Let $\phi_{\theta}[x,y]$ denote the network output, and define the interior residual
$r_{\theta}[x,y]\;=\;\phi_{\theta,xx}[x,y]+\phi_{\theta,yy}[x,y].$ (36)
We minimize the following objective over the interior sampling set $\mathcal{S}_{\Omega}$ and the four boundary sampling sets $\mathcal{S}_{y=0}$, $\mathcal{S}_{y=1}$, $\mathcal{S}_{x=0}$, $\mathcal{S}_{x=1}$:
$\begin{split}\mathcal{J}(\theta)&=\lambda_{\Omega}\,\frac{1}{|\mathcal{S}_{\Omega}|}\sum_{[x,y]\in\mathcal{S}_{\Omega}}\bigl|r_{\theta}[x,y]\bigr|^{2}\;+\;\lambda_{D0}\,\frac{1}{|\mathcal{S}_{y=0}|}\sum_{x\in\mathcal{S}_{y=0}}\bigl|\phi_{\theta}[x,0]\bigr|^{2}\\[4.0pt] &\quad+\;\lambda_{D1}\,\frac{1}{|\mathcal{S}_{y=1}|}\sum_{x\in\mathcal{S}_{y=1}}\bigl|\phi_{\theta}[x,1]-1\bigr|^{2}\;+\;\lambda_{N0}\,\frac{1}{|\mathcal{S}_{x=0}|}\sum_{y\in\mathcal{S}_{x=0}}\bigl|\partial_{x}\phi_{\theta}[0,y]\bigr|^{2}\;+\;\lambda_{N1}\,\frac{1}{|\mathcal{S}_{x=1}|}\sum_{y\in\mathcal{S}_{x=1}}\bigl|\partial_{x}\phi_{\theta}[1,y]\bigr|^{2}.\end{split}$ (37)
We obtain first- and second-order derivatives via automatic differentiation. We match the implementation’s sampling sizes: we draw $1000$ interior points and $1000$ points on each of the four boundaries.
In Fig. 5 we present the comparison. For xLSTM-PINN, the predicted and reference iso-bands almost perfectly overlap; the absolute-error map stays deep blue over the entire domain, with only fine and uniform vertical micro-striations whose maximum magnitude is about $3.8\times 10^{-4}$, essentially at the level of numerical noise. For the baseline PINN, the absolute error forms a smooth and wide warm-colored curved band of order $10^{-2}$ across the domain and exhibits low-order shape bias along the $y$-direction. We attribute the gap to the mechanism established in Sec 2: gated memory and residual micro-steps in xLSTM raise effective convergence weights for high-frequency directions, markedly suppress spectral bias, and therefore drive the system error down to the noise level under the same budget.
In Tab. 2 we report four metrics: MSE drops from $6.04\times 10^{-5}$ to $1.47\times 10^{-8}$; RMSE drops from $7.77\times 10^{-3}$ to $1.21\times 10^{-4}$; MAE drops from $7.75\times 10^{-3}$ to $9.90\times 10^{-5}$; MaxAE drops from $8.81\times 10^{-3}$ to $3.82\times 10^{-4}$. These reductions align with the spatial error patterns in Fig. 5. The training loss histories, presented in Fig. 6, further validate the improved convergence characteristics of xLSTM-PINN.
Table 2: Error metrics for the 2D Laplace mixed Dirichlet–Neumann case: MSE, RMSE, MAE, and MaxAE of xLSTM–PINN vs PINN; 1,000 interior samples, 1,000 per boundary side.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 1.47e-08 1.21e-04 9.90e-05 3.82e-04
PINN 6.04e-05 7.77e-03 7.75e-03 8.81e-03
Figure 6: Training loss histories for the 2D Laplace mixed-boundary case under matched sampling and optimization. (a) xLSTM–PINN reaches a low-loss regime faster and sustains a lower loss floor; (b) PINN on a logarithmic axis highlights slower late-phase decay and larger oscillations.
3.3 Non-dimensional Steady-State Heat Conduction in a Circular Silicon Plate with Convective Boundary
Figure 7: Steady heat conduction in a disk (uniform volumetric source ++ Robin convective boundary): using the FEM field as the validation set, we compare xLSTM-PINN and PINN on solution, PDE prediction, and absolute error. Domain $\Omega=\{x^{2}+y^{2}\leq 1\}$; PDE $\theta_{xx}+\theta_{yy}+1=0$; boundary $\partial_{n}\theta=\text{Bi}\,\theta$. (a) xLSTM-PINN: concentric isotherms with a thin cool error ring near the rim. (b) PINN: a wider warm error ring penetrating inward.
We start from the dimensional heat–transfer model. We assume the silicon wafer is thin enough to justify a two–dimensional approximation, and we define the physical temperature field $T(X,Y)$ on the disk
$D_{R}=\{(X,Y):\,X^{2}+Y^{2}\leq R^{2}\}.$ (38)
Using the ambient temperature $T_{\infty}$ as the reference, we set the temperature rise $\Delta T=T-T_{\infty}$. With constant conductivity $k$ and a uniform volumetric heat source $Q\,[\mathrm{W/m}^{3}]$, we write the governing equation and the Newton cooling boundary:
$k\!\left(\frac{\partial^{2}\Delta T}{\partial X^{2}}+\frac{\partial^{2}\Delta T}{\partial Y^{2}}\right)+Q=0\;\;\text{in }D_{R},\qquad-k\,\frac{\partial\Delta T}{\partial n}=h\,\Delta T\;\;\text{on }\partial D_{R}.$ (39)
We nondimensionalize the model. We choose the characteristic length $R$ and the temperature scale $T_{\mathrm{ref}}=\dfrac{QR^{2}}{k}$, and define
$x=\frac{X}{R},\qquad y=\frac{Y}{R},\qquad\theta(x,y)=\frac{\Delta T(X,Y)}{T_{\mathrm{ref}}},\qquad\Omega=\{(x,y):\,x^{2}+y^{2}\leq 1\}.$
Using $\partial_{X}=(1/R)\partial_{x}$ and $\partial_{Y}=(1/R)\partial_{y}$, we obtain the dimensionless model
$\theta_{xx}+\theta_{yy}+1=0\;\;\text{in }\Omega,\qquad-\partial_{n}\theta=\mathrm{Bi}\,\theta\;\;\text{on }\partial\Omega,$ (40)
where the Biot number $\mathrm{Bi}=hR/k$. We adopt the same physical parameters and boundary settings as the finite–element benchmark in [tao2025lnn], and we use its solution of FEM as a pointwise validation set.
Let $\theta_{\theta}(x,y)$ denote the network approximation. We define the interior residual
$r_{\theta}(x,y)=\theta_{\theta,xx}(x,y)+\theta_{\theta,yy}(x,y)+1,$
and the boundary residual
$b_{\theta}(x,y)=-\nabla\theta_{\theta}(x,y)\!\cdot n(x,y)-\mathrm{Bi}\,\theta_{\theta}(x,y).$
We minimize a weighted empirical $\ell_{2}$ objective over interior samples $\mathcal{S}_{\Omega}$ and boundary samples $\mathcal{S}_{\partial\Omega}$:
$\mathcal{J}(\theta)=\lambda_{\Omega}\frac{1}{|\mathcal{S}_{\Omega}|}\sum_{(x,y)\in\mathcal{S}_{\Omega}}\!\bigl|r_{\theta}(x,y)\bigr|^{2}\;+\;\lambda_{\partial\Omega}\frac{1}{|\mathcal{S}_{\partial\Omega}|}\sum_{(x,y)\in\mathcal{S}_{\partial\Omega}}\!\bigl|b_{\theta}(x,y)\bigr|^{2}.$ (41)
Automatic differentiation supplies first– and second–order derivatives. We match the implementation’s sampling sizes: $|\mathcal{S}_{\Omega}|=3000$ in the interior and $|\mathcal{S}_{\partial\Omega}|=500$ on the boundary.
In Fig. 7 we compare predictions with the FE validation set. xLSTM-PINN produces concentric contour lines that align with the analytic parabolic profile; the absolute–error map stays cold across the disk and forms only an ultrathin ring near the rim, with magnitude on the order of $10^{-4}$. The baseline PINN develops a wider and brighter boundary ring that diffuses inward into slowly undulating curved errors, at magnitudes one to two orders larger. We attribute this gap to the frequency–domain mechanism established in Sec 2: gated memory and residual micro–steps raise effective convergence weights along high–frequency directions, markedly suppress spectral bias, and therefore satisfy the Poisson source and the Robin flux more accurately under the same training budget.
In Tab. 3 we report four metrics consistent with this error structure: xLSTM-PINN reduces MSE from $2.86\times 10^{-7}$ to $9.66\times 10^{-9}$, RMSE from $5.35\times 10^{-4}$ to $9.83\times 10^{-5}$, MAE from $5.26\times 10^{-4}$ to $7.87\times 10^{-5}$, and MaxAE from $7.46\times 10^{-4}$ to $4.03\times 10^{-4}$. The training loss histories, presented in Fig. 8, further validate the improved convergence characteristics of xLSTM-PINN.
Table 3: Disk steady conduction — error metrics with uniform volumetric source and Robin convective boundary: MSE, RMSE, MAE, and MaxAE of xLSTM–PINN vs PINN against the FEM validation set from [tao2025lnn]; 3,000 interior samples and 500 boundary samples
Model MSE RMSE MAE MaxAE
xLSTM–PINN 9.66e-09 9.83e-05 7.87e-05 4.03e-04
PINN 2.86e-07 5.35e-04 5.26e-04 7.46e-04
Figure 8: Training loss histories for the disk conduction case (reference only; matched sampling and optimization). (a) xLSTM–PINN reaches a low-loss regime faster and sustains a lower floor; (b) logarithmic view of the same curves highlighting faster late-phase decay and smaller oscillations.
3.4 Anisotropic Poisson–Beam Equation
Figure 9: Anisotropic Poisson-Beam equation comparison on $0,1]^{2}$ with $u_{xx}-u_{yyyy}=(2-x^{2})e^{-y}$; boundary conditions given in the text. (a) xLSTM-PINN: solution aligns with prediction, absolute-error map shows weak ripples and reduced high-frequency oscillations; (b) PINN: larger absolute errors with broad undulations and stripe patterns.
We solve on the unit square $\Omega=(0,1)^{2}$ the fourth–order mixed operator
$u_{xx}-u_{yyyy}=f(x,y),\qquad f(x,y)=(2-x^{2})\,\mathrm{e}^{-y},$ (42)
and impose the boundary conditions
$u(x,0)=x^{2},\quad u_{yy}(x,0)=x^{2},\quad u(x,1)=\frac{x^{2}}{\mathrm{e}},\quad u_{yy}(x,1)=\frac{x^{2}}{\mathrm{e}},$ (43)
$u(0,y)=0,\quad u(1,y)=\mathrm{e}^{-y}.$ (44)
We validate with the analytic solution
$u^{*}(x,y)=x^{2}\,\mathrm{e}^{-y}.$ (45)
We specify sampling sets and their cardinalities. We draw the interior residual set $S_{r}\subset\Omega$ by i.i.d. uniform sampling with $|S_{r}|=N=1000$. On the boundary, we sample
$S_{y=0}^{(u)}=\{(x,0)\},\quad S_{y=1}^{(u)}=\{(x,1)\},\quad S_{y=0}^{(yy)}=\{(x,0)\},\quad S_{y=1}^{(yy)}=\{(x,1)\},\quad S_{x=0}=\{(0,y)\},\quad S_{x=1}=\{(1,y)\},$ (46)
and set the same size for each set, $|S_{\bullet}|=N_{\partial}=1000$. In this way, we enforce the PDE residual in the interior, match $u$ and $u_{yy}$ on the two horizontal edges, and impose Dirichlet data on the two vertical edges.
Using automatic differentiation, we obtain $u_{\theta,xx}$, $u_{\theta,yyyy}$, and $u_{\theta,yy}$ for the network output $u_{\theta}$. We minimize the weighted squared objective
$\begin{split}\mathcal{J}(\theta)&=\lambda_{r}\,\frac{1}{|S_{r}|}\sum_{(x,y)\in S_{r}}\Big|\,u_{\theta,xx}(x,y)-u_{\theta,yyyy}(x,y)-f(x,y)\,\Big|^{2}\\ &\quad+\lambda_{0}^{(u)}\,\frac{1}{|S_{y=0}^{(u)}|}\sum_{x\in S_{y=0}^{(u)}}\Big|\,u_{\theta}(x,0)-x^{2}\,\Big|^{2}\;+\;\lambda_{1}^{(u)}\,\frac{1}{|S_{y=1}^{(u)}|}\sum_{x\in S_{y=1}^{(u)}}\Big|\,u_{\theta}(x,1)-\tfrac{x^{2}}{\mathrm{e}}\,\Big|^{2}\\ &\quad+\lambda_{0}^{(yy)}\,\frac{1}{|S_{y=0}^{(yy)}|}\sum_{x\in S_{y=0}^{(yy)}}\Big|\,u_{\theta,yy}(x,0)-x^{2}\,\Big|^{2}\;+\;\lambda_{1}^{(yy)}\,\frac{1}{|S_{y=1}^{(yy)}|}\sum_{x\in S_{y=1}^{(yy)}}\Big|\,u_{\theta,yy}(x,1)-\tfrac{x^{2}}{\mathrm{e}}\,\Big|^{2}\\ &\quad+\lambda_{L}\,\frac{1}{|S_{x=0}|}\sum_{y\in S_{x=0}}\Big|\,u_{\theta}(0,y)\,\Big|^{2}\;+\;\lambda_{R}\,\frac{1}{|S_{x=1}|}\sum_{y\in S_{x=1}}\Big|\,u_{\theta}(1,y)-\mathrm{e}^{-y}\,\Big|^{2}.\end{split}$ (47)
We generate a validation set from $u^{*}$ and compute MSE, RMSE, MAE, and MaxAE on a dense grid; Tab. 4 summarizes these metrics. The visual comparison in Fig. 9 shows that xLSTM-PINN produces a cleaner upper-right corner and transition bands near boundaries, achieves markedly lower peak amplitudes than the baseline PINN, and attenuates high-frequency error ripples. These observations confirm that we effectively suppress spectral bias and improve overall accuracy for this stiff, high-order constrained problem. We keep $N$ and all boundary set sizes identical to the values stated above.
Table 4: Error metrics for the anisotropic Poisson–Beam case: xLSTM-PINN vs PINN evaluated against the analytic solution $u^{*}(x,y)=x^{2}e^{-y}$ ; reported metrics are MSE, RMSE, MAE, and MaxAE. Interior set $|S_{r}|=1000$; each boundary set has 1000 samples.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 1.93e-06 1.39e-03 1.09e-03 5.32e-03
PINN 5.94e-05 7.71e-03 7.66e-03 9.93e-03
Figure 10: Training loss histories for the anisotropic Poisson–Beam case (reference only; matched sampling and optimization). (a) xLSTM–PINN quickly clears the initial peak induced by fourth-order stiffness and then decays with smaller oscillations, sustaining a lower loss floor; (b) PINN decays more slowly with higher, denser late-phase spikes, indicating harder convergence on high-frequency components.
4 Conclusion
[…]
Declaration of competing interest
The authors declared that they have no conflicts of interest to this work.
Acknowledgment
This work is supported by the developing Project of Science and Technology of Jilin Province (20240402042GH).
Data availability
Data will be made available on request.