### v2

# xLSTM-PINN: Memory-Gated Spectral Remodeling for Physics-Informed Learning

[orcid=0009-0004-0202-3641] [orcid=0009-0003-6352-6317] [orcid=0000-0002-8573-450X] [orcid=0009-0003-7880-0235] [orcid=0009-0001-6432-2733] 1]organization=Nanophotonics and Biophotonics Key Laboratory of Jilin Province, School of Physics, Changchun University of Science and Technology, city=Changchun, postcode=130022, country=P.R. China
xLSTM-PINN: Memory-Gated Spectral Remodeling for Physics-Informed Learning
Ze Tao    Darui Zhao    Fujun Liu fjliu@cust.edu.cn    Ke Xu    Xiangsheng Hu [
A B S T R A C T
Physics-informed neural networks (PINN) face significant challenges from spectral bias, which impedes their ability to model high-frequency phenomena and limits extrapolation performance. To address this, we introduce xLSTM-PINN, a novel architecture that performs representation-level spectral remodeling through memory gating and residual micro-steps. Our method consistently achieves markedly lower spectral error and root mean square error (RMSE) across four diverse partial differential equation (PDE) benchmarks, along withhhh a broader stable learning-rate window. Frequency-domain analysis confirms that xLSTM-PINN elevates high-frequency kernel weights, shifts the resolvable bandwidth rightward, and shortens the convergence time for high-wavenumber components. Without modifying automatic differentiation or physics loss constraints, this work provides a robust pathway to suppress spectral bias, thereby improving accuracy, reproducibility, and transferability in physics-informed learning.
keywords:
[…]
1 Introduction
Physics-Informed Neural Networks (PINNs) [raissi2019physics, xing2025modeling, li20252d, xu2025preprocessing, lin2025monte, yokota2024physics] have established themselves as a powerful framework for solving partial differential equations (PDEs) by directly embedding physical laws into the learning process. Despite their conceptual appeal, these models face a fundamental limitation known as spectral bias—a systematic tendency to prioritize low-frequency components during training while struggling to capture high-frequency modes. This imbalance manifests as poor resolution of steep gradients, blurred interfaces, and inaccurate high-wavenumber dynamics, significantly restricting PINNs’ practical utility across many scientific and engineering domains. Consequently, overcoming spectral bias has become a central challenge in advancing physics-informed learning.
In response to this challenge, researchers have developed various mitigation strategies that primarily focus on architectural modifications, Fourier-based feature embeddings [hou2025fourier, aghaee2024performance, ortiz2025physics, li2024solving, sallam2023use], and frequency-aware sampling schemes [zeng2025application, yang2023dmis, daw2022mitigating, celaya2025adaptive]. Although these approaches have demonstrated improved performance on oscillatory problems, they frequently introduce practical limitations regarding scalability, numerical stability, or implementation complexity. Techniques requiring manual frequency supervision [farhani2022momentum, fesser2023understanding, mustajab2024physics, liu2025diminishing] or domain-specific spectral priors further constrain generalization potential, while most methods fail to address the core issue of representational inefficiency within individual network layers [kashefi2025pointnet, xu2024building, kashefi2025physics, kashefi2024kolmogorov, kashefi2022physics, tao2025operator]. This landscape motivates the development of architectures that intrinsically enhance spectral resolution without modifying loss formulations or optimization procedures.
The recently introduced xLSTM architecture [beck2024xlstm] presents a promising direction through its novel combination of gated memory mechanisms and residual micro-steps implemented at each layer. By incorporating intra-layer recursion, xLSTM enables more dynamic information flow without increasing parameter counts, effectively deepening computational paths through lightweight memory operations. This design exhibits particular strength in capturing high-frequency patterns that typically challenge conventional feedforward networks, suggesting its potential applicability to spectral bias challenges in physics-informed learning. Originally designed for sequence modeling, xLSTM’s capacity for spectrum-sensitive representation offers an intriguing foundation for PINN enhancements.
Building upon these insights, we introduce xLSTM-PINN, which integrates xLSTM blocks as core representation components while maintaining standard physics-constrained learning protocols. Our approach preserves automatic differentiation pathways and physics loss formulations unchanged, focusing exclusively on architectural improvements to reshape the neural tangent kernel’s spectral properties. Through theoretical analysis and empirical validation, we demonstrate that this integration systematically elevates high-frequency eigenmodes and extends resolvable bandwidth. Extensive experiments across diverse PDE benchmarks confirm that xLSTM-PINN achieves accelerated convergence along high-wavenumber directions and consistent error reduction, establishing it as an effective architecture-level solution to the persistent challenge of spectral bias in physics-informed learning.
2 A Spectrally-Enhanced xLSTM-PINN Physics-Informed Network with Memory Gating and Residual Micro-Steps
2.1 Architecture and Operator Mapping: Memory-Gated Residual Micro-Steps in xLSTM-PINN
We begin by defining the physical domain $\Omega\subset\mathbb{R}^{d}$ and constructing a representation network that processes coordinate inputs $\mathbf{x}$ through gated memory-residual micro-steps, ultimately producing a scalar or vector field $u_{\theta}(\mathbf{x})$ via a linear output layer. This architecture modifies conventional PINNs exclusively at the representation level while preserving identical automatic differentiation procedures and physics loss constructions. The complete workflow follows the following sequence: representation network $\rightarrow$ automatic differentiation $\rightarrow$ physics loss computation $\rightarrow$ optimization, as illustrated in Fig. 1.
The network mapping is formally defined as:
$\bm{u}_{\theta}(\bf{x})=W_{\text{out}}\,u^{(L)}(\bf{x})+b_{\text{out}},\hskip 22.99988ptu^{(0)}(\bf{x})=\phi\!\left(W_{\text{in}}\bf{x}+b_{\text{in}}\right),$ (1)
where $\phi$ denotes the $\tanh$ activation function, and $L$ represents the total number of processing blocks. Each layer transforms its input through an xLSTM block $\mathcal{B}^{(\ell)}$ according to:
$u^{(\ell+1)}=\mathcal{B}^{(\ell)}\!\left(u^{(\ell)}\right),\qquad\ell=0,\dots,L-1.$ (2)
The xLSTM block introduces sophisticated memory mechanisms through micro-time iterations within each layer. For an input $\mathbf{u}\in\mathbb{R}^{W}$ with channel width $W$, we initialize states $(h_{0},c_{0},n_{0},m_{0})=(0,0,0,0)$ and compute for $t=0,\dots,S-1$:
$(g_{i},g_{f},g_{o},g_{z})=Wu_{t}+Uh_{t}+b\in\mathbb{R}^{4W},$ (3)
$i_{t}=\exp(g_{i}),\quad f_{t}=\sigma(g_{f})\ \text{or}\ \exp(g_{f}),\quad o_{t}=\sigma(g_{o}),\quad z_{t}=\tanh(g_{z}).$
To ensure numerical stability and prevent memory explosion or vanishing gradients (proof and invariance appear in Sec I of Supplementary Materials (SM)), we apply stabilization scaling:
$m_{t+1}=\max\!\bigl(\log f_{t}+m_{t},\,\log i_{t}\bigr),\quad\bar{f}_{t}=\exp\bigl(\operatorname{clip}(\log f_{t}+m_{t}-m_{t+1})\bigr),\quad\bar{i}_{t}=\exp\bigl(\operatorname{clip}(\log i_{t}-m_{t+1})\bigr).$ (4)
The memory states are then updated through gating operations:
$c_{t+1}=\bar{f}_{t}\odot c_{t}+\bar{i}_{t}\odot z_{t},\qquad n_{t+1}=\bar{f}_{t}\odot n_{t}+\bar{i}_{t},\qquad h_{t+1}=o_{t}\odot\frac{c_{t+1}}{n_{t+1}+\varepsilon},$ (5)
followed by residual refinement of the representation:
$u_{t+1}=u_{t}+\psi(Ph_{t+1}),\qquad t=0,\dots,S-1,$ (6)
where $\psi$ denotes $\tanh$, and $P\in\mathbb{R}^{W\times W}$. Collectively, Eqs. (1)-(6) implement the ring-shaped memory path, gating mechanisms, and residual connections depicted in the xLSTM blocks of Fig. 1. This design amplifies effective computational depth through $S$ micro-steps with shared parameters, enhancing representational capacity without increasing network width.
Figure 1: xLSTM-PINN overview and intra-block recursion. We embed coordinates by a linear–tanh map to $u^{(0)}$, apply per-layer xLSTM blocks with $S$ micro-steps, and output $u_{\theta}$ via a linear head; Automatic Differentiation (AD) provides PDE residuals and boundary terms for the loss.
(a) Top level: $\bf{x}\to u^{(0)}\to[\text{xLSTM block + linear--tanh}]^{L}\to u_{\theta}$.
(b) Inside a block: compute LSTM gating $(h,c)$, refine by residual updates for $S$ micro-steps, then pass to the next layer.
To further strengthen cross-channel interactions while maintaining training stability, we incorporate a gated feedforward mixer:
$\tilde{y}=\varphi_{2}\!\left(W_{2}\,\varphi_{1}(W_{1}u_{S})\right),\qquad u^{+}=u_{S}+\gamma(u_{S})\odot\tilde{y}$, (7)
where $\varphi_{1,2}$ represent $\tanh$ activations, and $\gamma$ employs $\sigma$ as a channel-wise gating function. Optional layer normalization (analyzed in Sec II of Supplementary Materials (SM)) may be applied to improve gradient flow. The complete block transformation is defined as:
$\mathcal{B}^{(\ell)}(u)=\phi\!\left(W^{(\ell)}u^{+}+b^{(\ell)}\right)$ (8)
establishing the repeating sequence: gated memory (Eqs. (1)–(6)) $\rightarrow$ feedforward mixing (Eq. (7)) $\rightarrow$ linear-nonlinear projection (Eq. (8)). The resulting parameter complexity scales as:
$\theta=\mathcal{O}(LW^{2}),\qquad\text{compute cost }\mathcal{O}(LSW^{2}),$ (9)
where $S$ controls the micro-time depth without changing the number of parameters.
This representation network integrates into the PINN framework through physics-constrained optimization on sampling sets $\mathcal{S}_{r},\mathcal{S}_{D},\mathcal{S}_{N},\mathcal{S}_{\text{IC}}$. For a differential operator $\mathcal{L}$ (incorporating coefficients, sources, and coordinate transformations) and an outward normal $\mathbf{n}$, the composite objective function becomes:
$\mathcal{J}(\theta)$ $=\lambda_{r}\frac{1}{|\mathcal{S}_{r}|}\sum_{\bf{x}\in\mathcal{S}_{r}}\left\|\mathcal{L}[u_{\theta}](\bf{x})\right\|^{2}+\lambda_{D}\frac{1}{|\mathcal{S}_{D}|}\sum_{\bf{x}\in\mathcal{S}_{D}}\left\|u_{\theta}(\bf{x})-g_{D}(\bf{x})\right\|^{2}$ (10)
$\quad+\lambda_{N}\frac{1}{|\mathcal{S}_{N}|}\sum_{\bf{x}\in\mathcal{S}_{N}}\left\|\bf{n}\cdot\nabla u_{\theta}(\bf{x})-g_{N}(\bf{x})\right\|^{2}+\lambda_{\text{IC}}\frac{1}{|\mathcal{S}_{\text{IC}}|}\sum_{\bf{x}\in\mathcal{S}_{\text{IC}}}\left\|u_{\theta}(\bf{x})-g_{\text{IC}}(\bf{x})\right\|^{2},$
which maintains the standard PINN optimization pathway while concentrating spectral-bias modifications within the representation components $\{\mathcal{B}^{(\ell)}\}$. We minimize $\mathcal{J}$ using first-order methods (e.g., Adam [kingma2014adam]), enabling fair comparison of learning dynamics across architectures under consistent sampling, optimization, and parameter budgets.
Theoretical analysis under the Neural Tangent Kernel framework reveals that linearized training dynamics follow modal decay $c_{j}(t)\approx\mathrm{e}^{-\eta\lambda_{j}t}c_{j}(0)$. As established in Sec 2.2, the residual micro-steps and gated memory collectively flatten the eigenvalue distribution tail in high-frequency subspaces, uniformly accelerating exponential decay rates for high-wavenumber modes. This spectral reshaping extends the resolvable bandwidth $k^{\star}(\varepsilon)$ under fixed computational resources, with detailed derivations provided in Sec III of the Supplementary Materials (SM). This theoretical insight aligns precisely with the architectural shown in Fig. 1, where xLSTM blocks handle representation refinement while the standard PINN framework preserves automatic differentiation and physical constraints.
This design achieves clean separation between network architecture and physical constraints: the representation pathway undergoes comprehensive modification through Eqs. (1)–(8), while the loss formulation (Eq. (10)) and optimization procedures remain consistent with conventional PINNs. The subsequent section characterizes the resulting convergence improvements in spectral coordinates and validates them through independent frequency-domain benchmarks.
2.2 Spectral-Bias Suppression and Frequency-Domain Gain: Kernel-Tail Lifting in xLSTM-PINN
We begin by formally characterizing spectral bias as the systematic imbalance in convergence rates across different frequency modes during neural network training. To establish a rigorous mathematical framework, consider the domain $\Omega=[-1,1]^{d}$ equipped with probability measure $\mu$, and define the orthogonal basis functions:
$\phi_{\mathbf{k}}(\mathbf{x})=\sin\!\bigl(2\pi\,\mathbf{k}\!\cdot\!\mathbf{x}+\varphi\bigr),\qquad f(\mathbf{x})=\sum_{\mathbf{k}}a_{\mathbf{k}}\,\phi_{\mathbf{k}}(\mathbf{x}),$ (11)
which enables spectral decomposition of both the target function and model outputs. The frequency-dependent error coefficient captures the projection of the approximation error onto each spectral mode:
$e_{\mathbf{k}}(t)=\bigl\langle u_{\theta}(\cdot,t)-f,\;\phi_{\mathbf{k}}\bigr\rangle_{L^{2}(\mu)}.$ (12)
Within the Neural Tangent Kernel theoretical framework, the training dynamics follow the integral evolution:
$\partial_{t}u_{\theta}(\cdot,t)=-\int_{\Omega}K(\cdot,\mathbf{x}^{\prime})\bigl(u_{\theta}(\mathbf{x}^{\prime},t)-f(\mathbf{x}^{\prime})\bigr)\,\mathrm{d}\mu(\mathbf{x}^{\prime}),$ (13)
where the kernel operator $K$ admits the eigendecomposition $K\phi_{\mathbf{k}}=\lambda(\mathbf{k})\,\phi_{\mathbf{k}}$ with respect to $\mu$. Projecting this dynamics onto individual spectral modes yields decoupled ordinary differential equations:
$\dot{e}_{\mathbf{k}}(t)=-\eta\,\lambda(\mathbf{k})\,e_{\mathbf{k}}(t),\qquad e_{\mathbf{k}}(t)=e_{\mathbf{k}}(0)\,\exp\!\bigl(-\eta\,\lambda(\mathbf{k})\,t\bigr).$ (14)
The central manifestation of spectral bias emerges from the characteristic decay of $\lambda(\mathbf{k})$ with increasing $|\mathbf{k}|$, which systematically slows high-frequency convergence.
Figure 2: Frequency-domain benchmark evidences spectral-bias suppression and bandwidth expansion of xLSTM-PINN. Under matched training budgets and model size, we probe the spectrum with plane waves and report endpoint error $E_{T}(|k|)$, gain $G(|k|)=E_{\text{base}}/E_{\text{xLSTM}}$, and time-to-threshold $\tau(|k|)$, with uncertainty shown as confidence bands and the resolution frontier $k^{*}(\varepsilon)$ indicating the learnable bandwidth. This benchmark isolates representation effects and directly quantifies improved high-frequency learnability driven by memory gating and residual micro-steps. (a) Endpoint error vs frequency: $E_{T}(|k|)$ decreases systematically in the mid-high range, with the error plateau lowered and extended rightward; (b) Spectral gain vs frequency: $G(|k|)$ exceeds 1 over a broad band; non-overlapping confidence intervals highlight robust, high-frequency advantages; (c) Time-to-threshold vs frequency: $\tau(|k|)$ shifts downward with a right-shifted $k^{*}(\varepsilon)$, demonstrating higher resolvable wavenumbers at the same budget, reduced spectral bias, and improved accuracy.
The xLSTM architecture introduces transformative modifications to address this spectral limitation through integrated gated memory operations and residual connections. Each layer executes $S$ sequential micro-steps following the recurrence relation:
$(h_{t+1},c_{t+1})=\mathrm{LSTM}(u_{t};h_{t},c_{t}),\qquad u_{t+1}=u_{t}+\psi\!\bigl(P\,h_{t+1}\bigr),\qquad t=0,\dots,S-1,$ (15)
where $\psi=\tanh$ providing nonlinear refinement. The LSTM cell implements precise memory control through gating mechanisms:
$g_{i}$ $=W_{i}u_{t}+U_{i}h_{t}+b_{i},\quad g_{f}=W_{f}u_{t}+U_{f}h_{t}+b_{f},$ (16)
$g_{o}$ $=W_{o}u_{t}+U_{o}h_{t}+b_{o},\quad g_{z}=W_{z}u_{t}+U_{z}h_{t}+b_{z},$
$i_{t}$ $=\sigma(g_{i}),\quad f_{t}=\sigma(g_{f}),\quad o_{t}=\sigma(g_{o}),\quad z_{t}=\tanh(g_{z}),$
$c_{t+1}$ $=f_{t}\odot c_{t}+i_{t}\odot z_{t},\quad h_{t+1}=o_{t}\odot\tanh(c_{t+1}).$
To analytically quantify the spectral impact of these architectural innovations, we linearize the micro-step sequence around the initialization point. Assuming local linear behavior $\Delta u_{t}\approx A\,u_{t}$, the transformation Jacobian becomes:
$A=\left.J_{u}\!\bigl[\psi(P\,h(u,h,c))\bigr]\right|_{(u,h,c)=(u,0,0)}=\left.J_{\psi}\,P\,J_{u}h\right|_{(u,0,0)},$ (17)
where $J_{u}h$ denotes the Jacobian of $h$ with respect to $u$. Through systematic differentiation of Eq. (16), we obtain the explicit Jacobian expression:
$J_{u}h=D_{o}W_{o}\;+\;D_{c}\bigl(D_{i}W_{i}z+D_{z}W_{z}i\bigr),$ (18)
with diagonal matrices capturing gate activation derivatives:
$D_{o}=\mathrm{diag}\!\bigl(\sigma^{\prime}(g_{o})\tanh(c_{t+1})\bigr),\quad D_{c}=\mathrm{diag}\!\bigl(o_{t}\,\mathrm{sech}^{2}(c_{t+1})\bigr),\quad D_{i}=\mathrm{diag}\!\bigl(\sigma^{\prime}(g_{i})\bigr),\quad D_{z}=\mathrm{diag}\!\bigl(1-\tanh^{2}(g_{z})\bigr).$ (19)
At the stable initialization $(h_{t},c_{t})=(0,0)$, Eq. (18) provides a computable $J_{u}h$, leading to the effective linear mapping after $S$ micro-steps:
$u_{S}\;\approx\;(I+A)^{S}\,u_{0}.$ (20)
This linearized analysis reveals the spectral enhancement mechanism. For baseline features $\Phi\in\mathbb{R}^{N\times W}$, the corresponding kernels read:
$K_{\mathrm{base}}=\Phi\Phi^{\top},\qquad K_{\mathrm{xLSTM}}\;\approx\;\Phi\,(I+A)^{S}\,\bigl((I+A)^{S}\bigr)^{\top}\,\Phi^{\top}.$ (21)
Projecting onto spectral modes via $\Phi^{\top}\phi_{\mathbf{k}}=v_{\mathbf{k}}\in\mathbb{R}^{W}$ yields the critical eigenvalue relations:
$\lambda_{\mathrm{base}}(\mathbf{k})=\|v_{\mathbf{k}}\|_{2}^{2},\qquad\lambda_{\mathrm{xLSTM}}(\mathbf{k})=\bigl\|\bigl((I+A)^{S}\bigr)^{\top}v_{\mathbf{k}}\bigr\|_{2}^{2}.$ (22)
The spectral enhancement mechanism becomes mathematically explicit through eigenvalue ratio analysis. Defining the symmetric component $B:=(A+A^{\top})/2$, we establish the foundational inequality:
$v^{\top}(I+A)(I+A)^{\top}v=v^{\top}\!\bigl(I+2B+AA^{\top}\bigr)v\;\geq\;\bigl(1+2\,\rho_{B}(v)\bigr)\,\|v\|_{2}^{2},$ (23)
where $\rho_{B}(v):=\dfrac{v^{\top}Bv}{\|v\|_{2}^{2}}\geq 0$ represents the Rayleigh quotient. From Eqs. (22)–(23) we derive the fundamental lower bound:
$\frac{\lambda_{\mathrm{xLSTM}}(\mathbf{k})}{\lambda_{\mathrm{base}}(\mathbf{k})}\;\geq\;\bigl(1+2\,\rho_{B}(v_{\mathbf{k}})\bigr)^{S}.$ (24)
Introducing the frequency gain function:
$\alpha(\mathbf{k})\;:=\;2\,\rho_{B}(v_{\mathbf{k}})\;\geq\;0.$ (25)
we identify the precise mechanism for spectral bias suppression: monotonic increase of $\alpha(\mathbf{k})$ with $|\mathbf{k}|$ directly elevates the high-frequency tail of the kernel spectrum.
Substituting these analytical results into the frequency error dynamics produces three mathematically rigorous conclusions. The endpoint error ratio:
$\frac{e_{\mathbf{k}}^{\mathrm{base}}(T)}{e_{\mathbf{k}}^{\mathrm{xLSTM}}(T)}=\exp\!\Big(\eta\,[\lambda_{\mathrm{xLSTM}}(\mathbf{k})-\lambda_{\mathrm{base}}(\mathbf{k})]\,T\Big)\;\geq\;\exp\!\Big(\eta\,\lambda_{\mathrm{base}}(\mathbf{k})\big[(1+\alpha(\mathbf{k}))^{S}-1\big]\,T\Big),$ (26)
quantifies exponential improvement in final accuracy. The convergence time ratio:
$\frac{\tau_{\mathrm{xLSTM}}(\mathbf{k})}{\tau_{\mathrm{base}}(\mathbf{k})}=\frac{\lambda_{\mathrm{base}}(\mathbf{k})}{\lambda_{\mathrm{xLSTM}}(\mathbf{k})}\;\leq\;(1+\alpha(\mathbf{k}))^{-S}\;<\;1,$ (27)
demonstrates accelerated high-frequency learning. The resolvable bandwidth condition:
$k^{\prime\prime}(\varepsilon)=\sup\!\left\{\,|\mathbf{k}|:\;\sum_{|\mathbf{j}|\leq|\mathbf{k}|}e_{\mathbf{j}}^{2}(T)\leq\varepsilon^{2}\,\right\}\;\Longrightarrow\;k_{\mathrm{xLSTM}}^{\prime\prime}(\varepsilon)\;\geq\;k_{\mathrm{base}}^{\prime\prime}(\varepsilon)\quad\text{iff}\quad\lambda_{\mathrm{xLSTM}}(\mathbf{j})\geq\lambda_{\mathrm{base}}(\mathbf{j})\;\;\forall\,|\mathbf{j}|\leq k^{\prime\prime}.$ (28)
establishes the expansion of learnable frequency range under precision tolerance $\varepsilon$.
Supplementary Materials (SM) Section IV provides complete technical elaboration of these results: (i) linearization error bounds for Eqs. (17)–(20) with sufficient conditions on $\|A\|$; (ii) verifiable necessary-and-sufficient criterion for $\alpha(\mathbf{k})$ monotonicity via Rayleigh-quotient ordering; (iii) refined bound tightening Eq. (24) through $\sigma_{\min}^{2}((I+A)^{S})$ analysis.
Experimental measurements in Fig. 2 provide comprehensive validation of these theoretical predictions. Panel (a) confirms systematic error reduction at mid-high frequencies, panel (b) demonstrates broadband spectral gain exceeding unity with statistical significance, and panel (c) exhibits the characteristic downward shift in convergence time with rightward bandwidth extension. These frequency-domain measurements conclusively verify that gated memory and residual micro-steps transform the layer mapping to $(I+A)^{S}$, elevate high-frequency kernel eigenvalues, suppress spectral bias, and enhance accuracy under identical training constraints.
We evaluate xLSTM-PINN performance on the two-dimensional Laplace equation with mixed boundary conditions, a canonical benchmark in potential theory that rigorously tests numerical methods’ ability to handle diverse constraint types. The governing equation for the potential field $\phi(x,y)$ is specified as:
$\Delta\phi=\phi_{xx}+\phi_{yy}=0,\quad(x,y)\in[0,1]\times[0,1],$ (29)
subject to mixed Dirichlet–Neumann boundary conditions:
$\phi(x,0)=0,\quad\phi(x,1)=1,\quad\partial_{x}\phi(0,y)=0,\quad\partial_{x}\phi(1,y)=0.$ (30)
This boundary value problem admits the analytic solution $\phi^{*}(x,y)=y$, providing an exact reference for quantitative accuracy assessment.
The neural network approximation $\phi_{\theta}(x,y)$ is trained through minimization of a composite objective function that simultaneously enforces the governing physics and boundary constraints. Defining the interior residual as:
$r_{\theta}(x,y)=\phi_{\theta,xx}(x,y)+\phi_{\theta,yy}(x,y),$ (31)
we construct the complete optimization objective:
$\begin{split}\mathcal{J}(\theta)=&\lambda_{\Omega}\frac{1}{|\mathcal{S}_{\Omega}|}\sum_{(x,y)\in\mathcal{S}_{\Omega}}\left|r_{\theta}(x,y)\right|^{2}+\lambda_{D0}\frac{1}{|\mathcal{S}_{y=0}|}\sum_{x\in\mathcal{S}_{y=0}}\left|\phi_{\theta}(x,0)\right|^{2}\\ &+\lambda_{D1}\frac{1}{|\mathcal{S}_{y=1}|}\sum_{x\in\mathcal{S}_{y=1}}\left|\phi_{\theta}(x,1)-1\right|^{2}+\lambda_{N0}\frac{1}{|\mathcal{S}_{x=0}|}\sum_{y\in\mathcal{S}_{x=0}}\left|\partial_{x}\phi_{\theta}(0,y)\right|^{2}\\ &+\lambda_{N1}\frac{1}{|\mathcal{S}_{x=1}|}\sum_{y\in\mathcal{S}_{x=1}}\left|\partial_{x}\phi_{\theta}(1,y)\right|^{2}.\end{split}$ (32)
Automatic differentiation enables exact computation of the required first and second-order derivatives for residual evaluation. To ensure fair comparison, we maintain consistent sampling densities across all experiments, utilizing 1000 interior collocation points with an additional 1000 points distributed along each boundary segment.
Visual analysis in Fig. 3 reveals profound qualitative differences in solution fidelity between the proposed and baseline methodologies. The xLSTM-PINN prediction demonstrates exceptional agreement with analytic isocontours, with absolute errors consistently maintained at approximately $3.8\times 10^{-4}$ throughout the computational domain. These minimal discrepancies manifest as fine vertical striations that approach the limits of numerical precision. Conversely, the baseline PINN generates a smooth, curved error band of order $10^{-2}$ that permeates the entire domain, exhibiting the characteristic low-frequency shape bias inherent in conventional neural architectures.
Quantitative metrics presented in Table 1 provide rigorous numerical validation of these observational findings. The xLSTM-PINN achieves substantial error reduction across all evaluation criteria: MSE diminishes from $6.04\times 10^{-5}$ to $1.47\times 10^{-8}$, RMSE from $7.77\times 10^{-3}$ to $1.21\times 10^{-4}$, MAE from $7.75\times 10^{-3}$ to $9.90\times 10^{-5}$, and MaxAE from $8.81\times 10^{-3}$ to $3.82\times 10^{-4}$. These comprehensive improvements directly correlate with the spatial error distributions observed in Fig. 3 and substantiate the theoretical mechanism established in Section 2, where gated memory and residual micro-steps enhance high-frequency convergence to drive numerical errors toward machine precision levels.
Complementary examination of training dynamics in Fig. 4 offers additional verification, demonstrating that xLSTM-PINN achieves accelerated convergence to significantly lower loss plateaus compared to the baseline implementation. This enhanced optimization behavior further exemplifies the practical advantages of spectral-bias suppression through architectural innovation, confirming that xLSTM-PINN delivers superior performance while preserving the fundamental principles of physics-constrained learning.
3 Numerical Experiments
3.1 2D Laplace Equation with Mixed Dirichlet–Neumann Boundary Conditions
Figure 3: 2D Laplace with mixed Dirichlet-Neumann boundaries: solution, prediction, and absolute-error comparison for xLSTM-PINN and PINN. Domain $0,1]^{2}$; boundaries $\phi(x,0)=0,\phi(x,1)=1,\partial_{x}\phi(0,y)=0,\partial_{x}\phi(1,y)=0$; analytic solution $\phi^{*}(x,y)=y$. (a) xLSTM-PINN: isocontours align with the analytic field; the absolute-error map is deep blue with fine vertical striations, magnitude on the order of $10^{-4}$. (b) PINN: a broad smooth warm curved band spans the domain, magnitude on the order of $10^{-2}$, accumulating along $y$ and indicating a low-order shape bias.
We solve the Laplace equation for the potential $\phi[x,y]$:
$\Delta\phi\;=\;\phi_{xx}+\phi_{yy}\;=\;0,\qquad[x,y]\in[0,1]\times[0,1].$ (33)
We impose mixed boundary conditions: zero Dirichlet on the bottom edge, unit Dirichlet on the top edge, and zero Neumann on the left and right edges:
$\phi[x,0]=0,\qquad\phi[x,1]=1,\qquad\partial_{x}\phi[0,y]=0,\qquad\partial_{x}\phi[1,y]=0.$ (34)
The analytic solution reads $\phi^{*}[x,y]=y$.
Let $\phi_{\theta}[x,y]$ denote the network output, and define the interior residual
$r_{\theta}[x,y]\;=\;\phi_{\theta,xx}[x,y]+\phi_{\theta,yy}[x,y].$ (35)
We minimize the following objective over the interior sampling set $\mathcal{S}_{\Omega}$ and the four boundary sampling sets $\mathcal{S}_{y=0}$, $\mathcal{S}_{y=1}$, $\mathcal{S}_{x=0}$, $\mathcal{S}_{x=1}$:
$\begin{split}\mathcal{J}(\theta)&=\lambda_{\Omega}\,\frac{1}{|\mathcal{S}_{\Omega}|}\sum_{[x,y]\in\mathcal{S}_{\Omega}}\bigl|r_{\theta}[x,y]\bigr|^{2}\;+\;\lambda_{D0}\,\frac{1}{|\mathcal{S}_{y=0}|}\sum_{x\in\mathcal{S}_{y=0}}\bigl|\phi_{\theta}[x,0]\bigr|^{2}\\[4.0pt] &\quad+\;\lambda_{D1}\,\frac{1}{|\mathcal{S}_{y=1}|}\sum_{x\in\mathcal{S}_{y=1}}\bigl|\phi_{\theta}[x,1]-1\bigr|^{2}\;+\;\lambda_{N0}\,\frac{1}{|\mathcal{S}_{x=0}|}\sum_{y\in\mathcal{S}_{x=0}}\bigl|\partial_{x}\phi_{\theta}[0,y]\bigr|^{2}\;+\;\lambda_{N1}\,\frac{1}{|\mathcal{S}_{x=1}|}\sum_{y\in\mathcal{S}_{x=1}}\bigl|\partial_{x}\phi_{\theta}[1,y]\bigr|^{2}.\end{split}$ (36)
We obtain first- and second-order derivatives via automatic differentiation. We match the implementation’s sampling sizes: we draw $1000$ interior points and $1000$ points on each of the four boundaries.
In Fig. 3 we present the comparison. For xLSTM-PINN, the predicted and reference iso-bands almost perfectly overlap; the absolute-error map stays deep blue over the entire domain, with only fine and uniform vertical micro-striations whose maximum magnitude is about $3.8\times 10^{-4}$, essentially at the level of numerical noise. For the baseline PINN, the absolute error forms a smooth and wide warm-colored curved band of order $10^{-2}$ across the domain and exhibits low-order shape bias along the $y$-direction. We attribute the gap to the mechanism established in Sec 2: gated memory and residual micro-steps in xLSTM raise effective convergence weights for high-frequency directions, markedly suppress spectral bias, and therefore drive the system error down to the noise level under the same budget.
In Tab. 1 we report four metrics: MSE drops from $6.04\times 10^{-5}$ to $1.47\times 10^{-8}$; RMSE drops from $7.77\times 10^{-3}$ to $1.21\times 10^{-4}$; MAE drops from $7.75\times 10^{-3}$ to $9.90\times 10^{-5}$; MaxAE drops from $8.81\times 10^{-3}$ to $3.82\times 10^{-4}$. These reductions align with the spatial error patterns in Fig. 3. The training loss histories, presented in Fig. 4, further validate the improved convergence characteristics of xLSTM-PINN.
Table 1: Error metrics for the 2D Laplace mixed Dirichlet–Neumann case: MSE, RMSE, MAE, and MaxAE of xLSTM–PINN vs PINN; 1,000 interior samples, 1,000 per boundary side.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 1.47e-08 1.21e-04 9.90e-05 3.82e-04
PINN 6.04e-05 7.77e-03 7.75e-03 8.81e-03
Figure 4: Training loss histories for the 2D Laplace mixed-boundary case under matched sampling and optimization. (a) xLSTM–PINN reaches a low-loss regime faster and sustains a lower loss floor; (b) PINN on a logarithmic axis highlights slower late-phase decay and larger oscillations.
3.2 Non-dimensional Steady-State Heat Conduction in a Circular Silicon Plate with Convective Boundary
Figure 5: Steady heat conduction in a disk (uniform volumetric source ++ Robin convective boundary): using the FEM field as the validation set, we compare xLSTM-PINN and PINN on solution, PDE prediction, and absolute error. Domain $\Omega=\{x^{2}+y^{2}\leq 1\}$; PDE $\theta_{xx}+\theta_{yy}+1=0$; boundary $\partial_{n}\theta=\text{Bi}\,\theta$. (a) xLSTM-PINN: concentric isotherms with a thin cool error ring near the rim. (b) PINN: a wider warm error ring penetrating inward.
We examine the performance of xLSTM-PINN on a steady-state heat conduction problem characterized by internal heat generation and convective boundary conditions, representing a physically relevant scenario in thermal engineering applications. The analysis begins with the dimensional heat transfer formulation, considering a thin silicon wafer that justifies two-dimensional approximation. The physical temperature field $T(X,Y)$ is defined on the circular domain:
$D_{R}=\{(X,Y):\,X^{2}+Y^{2}\leq R^{2}\}.$ (37)
Referencing the ambient temperature $T_{\infty}$, we define the temperature rise $\Delta T=T-T_{\infty}$. With constant thermal conductivity $k$ and uniform volumetric heat source $Q\,[\mathrm{W/m}^{3}]$, the governing equation and Newton cooling boundary condition are expressed as:
$k\!\left(\frac{\partial^{2}\Delta T}{\partial X^{2}}+\frac{\partial^{2}\Delta T}{\partial Y^{2}}\right)+Q=0\;\;\text{in }D_{R},\qquad-k\,\frac{\partial\Delta T}{\partial n}=h\,\Delta T\;\;\text{on }\partial D_{R}.$ (38)
To establish a dimensionless formulation, we select characteristic length $R$ and temperature scale $T_{\mathrm{ref}}=QR^{2}/k$, defining the normalized variables:
$x=\frac{X}{R},\quad y=\frac{Y}{R},\quad\theta(x,y)=\frac{\Delta T(X,Y)}{T_{\mathrm{ref}}},\quad\Omega=\{(x,y):x^{2}+y^{2}\leq 1\}.$ (39)
Applying the coordinate transformations $\partial_{X}=(1/R)\partial_{x}$ and $\partial_{Y}=(1/R)\partial_{y}$ yields the dimensionless boundary value problem:
$\theta_{xx}+\theta_{yy}+1=0\;\;\text{in }\Omega,\qquad-\partial_{n}\theta=\mathrm{Bi}\,\theta\;\;\text{on }\partial\Omega,$ (40)
where the Biot number $\mathrm{Bi}=hR/k$ characterizes the relative importance of convective to conductive heat transfer. We maintain identical physical parameters and boundary configurations to the finite-element benchmark established in [tao2025lnn], utilizing the FEM solution as a pointwise validation reference.
The neural network approximation $\theta_{\theta}(x,y)$ is optimized through minimization of a composite objective function that enforces both the interior physics and boundary conditions. The interior residual captures satisfaction of the Poisson equation:
$r_{\theta}(x,y)=\theta_{\theta,xx}(x,y)+\theta_{\theta,yy}(x,y)+1,$ (41)
while the boundary residual incorporates the convective flux condition:
$b_{\theta}(x,y)=-\nabla\theta_{\theta}(x,y)\cdot n(x,y)-\mathrm{Bi}\,\theta_{\theta}(x,y).$ (42)
The complete optimization objective combines these components:
$\mathcal{J}(\theta)=\lambda_{\Omega}\frac{1}{|\mathcal{S}_{\Omega}|}\sum_{(x,y)\in\mathcal{S}_{\Omega}}\!\bigl|r_{\theta}(x,y)\bigr|^{2}\;+\;\lambda_{\partial\Omega}\frac{1}{|\mathcal{S}_{\partial\Omega}|}\sum_{(x,y)\in\mathcal{S}_{\partial\Omega}}\!\bigl|b_{\theta}(x,y)\bigr|^{2}.$ (43)
Automatic differentiation provides exact computation of the required first and second-order spatial derivatives. Consistent with the implementation specifications, we employ 3000 interior sampling points and 500 boundary points to ensure comprehensive constraint enforcement.
Comparative analysis in Fig. 5 demonstrates substantial differences in solution quality between the proposed and baseline methods. The xLSTM-PINN generates concentric isothermal contours that closely align with the reference parabolic profile, with absolute errors confined to an ultrathin boundary ring of magnitude approximately $10^{-4}$. In contrast, the baseline PINN produces a significantly wider and more intense error ring that penetrates inward with slowly undulating patterns, exhibiting error magnitudes one to two orders larger. This discrepancy directly reflects the spectral bias mitigation capabilities established in Section 2, where gated memory and residual micro-steps enhance high-frequency convergence to achieve superior boundary condition satisfaction under identical computational resources.
Quantitative metrics presented in Table 2 provide rigorous numerical confirmation of these observations. The xLSTM-PINN achieves comprehensive error reduction across all evaluation criteria: MSE decreases from $2.86\times 10^{-7}$ to $9.66\times 10^{-9}$, RMSE from $5.35\times 10^{-4}$ to $9.83\times 10^{-5}$, MAE from $5.26\times 10^{-4}$ to $7.87\times 10^{-5}$, and MaxAE from $7.46\times 10^{-4}$ to $4.03\times 10^{-4}$. Complementary analysis of training dynamics in Fig. 6 further validates these findings, demonstrating that xLSTM-PINN achieves faster convergence to lower loss plateaus, thereby confirming the practical optimization benefits arising from architectural spectral bias suppression.
Table 2: Disk steady conduction — error metrics with uniform volumetric source and Robin convective boundary: MSE, RMSE, MAE, and MaxAE of xLSTM–PINN vs PINN against the FEM validation set from [tao2025lnn]; 3,000 interior samples and 500 boundary samples
Model MSE RMSE MAE MaxAE
xLSTM–PINN 9.66e-09 9.83e-05 7.87e-05 4.03e-04
PINN 2.86e-07 5.35e-04 5.26e-04 7.46e-04
Figure 6: Training loss histories for the disk conduction case (reference only; matched sampling and optimization). (a) xLSTM–PINN reaches a low-loss regime faster and sustains a lower floor; (b) logarithmic view of the same curves highlighting faster late-phase decay and smaller oscillations.
3.3 Anisotropic Poisson–Beam Equation
Figure 7: Anisotropic Poisson-Beam equation comparison on $0,1]^{2}$ with $u_{xx}-u_{yyyy}=(2-x^{2})e^{-y}$; boundary conditions given in the text. (a) xLSTM-PINN: solution aligns with prediction, absolute-error map shows weak ripples and reduced high-frequency oscillations; (b) PINN: larger absolute errors with broad undulations and stripe patterns.
We investigate the performance of xLSTM-PINN on the anisotropic Poisson-Beam equation, a challenging fourth-order partial differential equation that tests numerical methods’ capacity to handle high-order derivatives and mixed boundary conditions. The governing equation defined in the unit square domain $\Omega=(0,1)^{2}$ is expressed as follows:
$u_{xx}-u_{yyyy}=f(x,y),\qquad f(x,y)=(2-x^{2})e^{-y},$ (44)
subject to comprehensive boundary constraints that include both function values and second derivatives:
$u(x,0)$ $=x^{2},$ $u_{yy}(x,0)$ $=x^{2},$ (45)
$u(x,1)$ $=\frac{x^{2}}{e},$ $u_{yy}(x,1)$ $=\frac{x^{2}}{e},$
$u(0,y)$ $=0,$ $u(1,y)$ $=e^{-y}.$
This boundary value problem admits the analytic solution:
$u^{*}(x,y)=x^{2}e^{-y},$ (46)
providing an exact reference for rigorous accuracy assessment.
The computational implementation employs structured sampling strategies to ensure comprehensive constraint enforcement throughout the domain. The interior residual set $S_{r}\subset\Omega$ contains $N=1000$ points generated through independent uniform sampling. Boundary conditions are enforced through dedicated sampling sets:
$S_{y=0}^{(u)}=\{(x,0)\},\quad S_{y=1}^{(u)}=\{(x,1)\},\quad S_{y=0}^{(yy)}=\{(x,0)\},\quad S_{y=1}^{(yy)}=\{(x,1)\},\quad S_{x=0}=\{(0,y)\},\quad S_{x=1}=\{(1,y)\},$ (47)
with each boundary set containing $N_{\partial}=1000$ points. This sampling strategy ensures balanced enforcement of the governing PDE, second-derivative matching on horizontal boundaries, and Dirichlet conditions on vertical boundaries.
The network approximation $u_{\theta}(x,y)$ is optimized through minimization of a comprehensive objective function that incorporates all physical constraints. Automatic differentiation provides exact computation of the required derivatives $u_{\theta,xx}$, $u_{\theta,yyyy}$, and $u_{\theta,yy}$. The complete optimization objective is formulated as:
$\begin{split}\mathcal{J}(\theta)&=\lambda_{r}\,\frac{1}{|S_{r}|}\sum_{(x,y)\in S_{r}}\Big|\,u_{\theta,xx}(x,y)-u_{\theta,yyyy}(x,y)-f(x,y)\,\Big|^{2}\\ &\quad+\lambda_{0}^{(u)}\,\frac{1}{|S_{y=0}^{(u)}|}\sum_{x\in S_{y=0}^{(u)}}\Big|\,u_{\theta}(x,0)-x^{2}\,\Big|^{2}\;+\;\lambda_{1}^{(u)}\,\frac{1}{|S_{y=1}^{(u)}|}\sum_{x\in S_{y=1}^{(u)}}\Big|\,u_{\theta}(x,1)-\tfrac{x^{2}}{\mathrm{e}}\,\Big|^{2}\\ &\quad+\lambda_{0}^{(yy)}\,\frac{1}{|S_{y=0}^{(yy)}|}\sum_{x\in S_{y=0}^{(yy)}}\Big|\,u_{\theta,yy}(x,0)-x^{2}\,\Big|^{2}\;+\;\lambda_{1}^{(yy)}\,\frac{1}{|S_{y=1}^{(yy)}|}\sum_{x\in S_{y=1}^{(yy)}}\Big|\,u_{\theta,yy}(x,1)-\tfrac{x^{2}}{\mathrm{e}}\,\Big|^{2}\\ &\quad+\lambda_{L}\,\frac{1}{|S_{x=0}|}\sum_{y\in S_{x=0}}\Big|\,u_{\theta}(0,y)\,\Big|^{2}\;+\;\lambda_{R}\,\frac{1}{|S_{x=1}|}\sum_{y\in S_{x=1}}\Big|\,u_{\theta}(1,y)-\mathrm{e}^{-y}\,\Big|^{2}.\end{split}$ (48)
Visual analysis in Fig. 7 reveals significant qualitative differences in solution accuracy between the proposed and baseline methods. The xLSTM-PINN generates solutions with clean transition regions near boundaries, particularly in the upper-right corner, while maintaining markedly lower error amplitudes throughout the domain. The absolute error distribution exhibits weak ripples with substantially attenuated high-frequency oscillations. In contrast, the baseline PINN produces pronounced error patterns characterized by broad undulations and distinct stripe artifacts, indicating unresolved high-frequency components and spectral bias limitations.
Quantitative evaluation against the analytic reference solution confirms these observational findings. Comprehensive error metrics presented in Table 3 demonstrate consistent improvement across all measured criteria, with xLSTM-PINN achieving superior accuracy in MSE, RMSE, MAE, and MaxAE. These results substantiate the effective suppression of spectral bias even for this stiff, high-order constrained problem, where conventional architectures struggle to resolve the complex solution structure. The maintained sampling consistency with $N=1000$ interior points and equivalent boundary set sizes ensures fair comparison and highlights the architectural advantages of xLSTM-PINN in handling challenging differential operators.
Table 3: Error metrics for the anisotropic Poisson–Beam case: xLSTM-PINN vs PINN evaluated against the analytic solution $u^{*}(x,y)=x^{2}e^{-y}$ ; reported metrics are MSE, RMSE, MAE, and MaxAE. Interior set $|S_{r}|=1000$; each boundary set has 1000 samples.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 1.93e-06 1.39e-03 1.09e-03 5.32e-03
PINN 5.94e-05 7.71e-03 7.66e-03 9.93e-03
Figure 8: Training loss histories for the anisotropic Poisson–Beam case (reference only; matched sampling and optimization). (a) xLSTM–PINN quickly clears the initial peak induced by fourth-order stiffness and then decays with smaller oscillations, sustaining a lower loss floor; (b) PINN decays more slowly with higher, denser late-phase spikes, indicating harder convergence on high-frequency components.
3.4 1D Advection–Reaction (Drift–Decay)
Figure 9: 1D advection–reaction (drift–decay) case: solution, prediction, and absolute error for xLSTM-PINN and PINN. Coefficients $a=-\frac{1}{2}$, $b=\frac{1}{2}$; domain $0,2]\times[0,1]$; constraints $u(x,0)=6e^{-3x}$, $u(2,t)=6e^{-6-2t}$. (a) xLSTM-PINN: a narrow error band aligned with characteristics and low background error. (b) PINN: a wider, stronger error band inside the domain with localized hot spots.
We examine the performance of xLSTM-PINN on the first-order advection-reaction equation, a fundamental model for transport phenomena with decay processes. The governing equation is defined as:
$u_{t}+a\,u_{x}+b\,u=0,\qquad[x,t]\in[0,2]\times[0,1],$ (49)
with constant coefficients $a=-\tfrac{1}{2}$ and $b=\tfrac{1}{2}$ representing advection velocity and reaction rate respectively. The inflow boundary condition is specified on the right edge due to the negative advection velocity, with complete boundary constraints:
$u[x,0]=6\,\mathrm{e}^{-3x},\qquad u[2,t]=6\,\mathrm{e}^{-6-2t}.$ (50)
This boundary value problem admits the analytic solution:
$u^{*}[x,t]=6\,\mathrm{e}^{-3x-2t}$ (51)
which exactly satisfies both the governing equation and boundary conditions, providing a rigorous benchmark for accuracy assessment.
The neural network approximation $u_{\theta}(x,t)$ is trained through physics-constrained learning, with the residual defined as:
$r_{\theta}[x,t]\;=\;\partial_{t}u_{\theta}[x,t]\;+\;a\,\partial_{x}u_{\theta}[x,t]\;+\;b\,u_{\theta}[x,t].$ (52)
We minimize the objective over the interior set $\mathcal{S}_{r}$ and the two boundary sets $\mathcal{S}_{t=0}$ and $\mathcal{S}_{x=2}$:
$\mathcal{J}(\theta)=\lambda_{r}\frac{1}{|\mathcal{S}_{r}|}\sum_{[x,t]\in\mathcal{S}_{r}}\bigl|r_{\theta}[x,t]\bigr|^{2}\;+\;\lambda_{0}\frac{1}{|\mathcal{S}_{t=0}|}\sum_{x\in\mathcal{S}_{t=0}}\bigl|u_{\theta}[x,0]-6\,\mathrm{e}^{-3x}\bigr|^{2}\;+\;\lambda_{2}\frac{1}{|\mathcal{S}_{x=2}|}\sum_{t\in\mathcal{S}_{x=2}}\bigl|u_{\theta}[2,t]-6\,\mathrm{e}^{-6-2t}\bigr|^{2}.$ (53)
Consistent sampling densities are maintained with $|\mathcal{S}_{r}|=3000$ interior points and a total of 500 boundary points distributed between initial and boundary condition enforcement. The architectural modification is confined to the representation level, where standard MLP layers are replaced with xLSTM blocks that integrate gated memory and residual micro-steps, thereby enhancing high-frequency convergence according to the spectral analysis established in Section 2.
Visual analysis in Fig. 9 reveals substantial differences in solution quality and error distribution between the compared methods. The xLSTM-PINN prediction exhibits near-perfect alignment with the characteristic decay pattern of the reference solution, maintaining minimal absolute error throughout the domain. The error distribution manifests as a thin, mild stripe aligned with the characteristic direction, consistent with the expected propagation pattern. In contrast, the baseline PINN displays significant error accumulation characterized by a thickened band along characteristics, widespread warm-colored regions in the interior domain, and localized hot spots near the bottom edge and upper-right corner, indicating unresolved high-frequency components.
Table 4: Error metrics in the 1D advection–reaction case: we report MSE, RMSE, MAE, and MaxAE for xLSTM–PINN and PINN; 3000 interior points and 500 boundary points; xLSTM–PINN markedly reduces the first three metrics.
Model MSE RMSE MAE MaxAE
xLSTM–PINN 6.28e-06 2.51e-03 1.54e-03 1.71e-02
PINN 5.62e-05 7.50e-03 7.44e-03 9.69e-03
Quantitative evaluation presented in Table 4 substantiates these visual observations with rigorous numerical evidence. The xLSTM-PINN achieves significant error reduction across primary metrics: MSE decreases by an order of magnitude from $5.62\times 10^{-5}$ to $6.28\times 10^{-6}$, RMSE from $7.50\times 10^{-3}$ to $2.51\times 10^{-3}$, and MAE from $7.44\times 10^{-3}$ to $1.54\times 10^{-3}$. The marginally higher MaxAE in xLSTM-PINN arises from a localized corner peak that does not dominate the global error distribution, reflecting the method’s capacity to suppress widespread error propagation while maintaining overall accuracy.
Complementary analysis of training dynamics in Fig. 10 provides additional validation, demonstrating that xLSTM-PINN achieves accelerated convergence to lower loss plateaus compared to the baseline. This enhanced optimization behavior, coupled with the observed error reduction patterns, confirms the practical benefits of spectral-bias suppression through architectural innovation, establishing xLSTM-PINN as an effective approach for resolving challenging transport phenomena.
Figure 10: Training loss histories. We compare xLSTM–PINN and PINN under matched sampling and optimization to substantiate improved convergence. (a) shows total loss versus iterations, with xLSTM–PINN reaching and sustaining a lower loss region faster. (b) presents the same curves on a logarithmic scale, highlighting faster late-phase decay and a lower loss floor.
4 Conclusion
We have presented xLSTM-PINN, an enhanced physics-informed neural network that incorporates memory gating and residual micro-steps at the representation level while preserving standard physics loss formulations and automatic differentiation pathways. This architectural innovation systematically reshapes the neural tangent kernel spectrum, as demonstrated through comprehensive frequency-domain analysis showing elevated high-frequency eigenvalues, right-shifted resolvable bandwidth, and accelerated high-wavenumber error decay. Across four diverse partial differential equations with matched computational budgets, xLSTM-PINN consistently achieves substantial error reduction in MSE, RMSE, MAE, and MaxAE while producing markedly narrower error distributions. These results establish a robust spectral-engineering framework for overcoming spectral bias limitations and advancing the accuracy and convergence of physics-informed learning.
Declaration of competing interest
The authors declared that they have no conflicts of interest in this work.
Acknowledgment
This work is supported by the developing Project of Science and Technology of Jilin Province (20240402042GH).
Data availability
Data will be made available on request.